OpenStudy (anonymous):

Perform the indicated operation. Please show all of your work. 5m/n÷2/3

OpenStudy (anonymous):

dividing by two thirds same as multiplying by three halves \[\frac{5m}{n}\times \frac{3}{2}\] multiply straight across